# Dynamic asymptotic dimension growth for group actions and groupoids

### v1

@@ Abstract.
We introduce the notion of dynamic asymptotic dimension growth for actions of discrete groups on compact spaces, and more generally for locally compact étale groupoids with compact unit space. Using the work of Bartels, Lück, and Reich to bridge our notion for group actions with asymptotic dimension growth for groups, thus providing many concrete examples for our notion. Moreover, we show that the asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid, hence deducing that the coarse groupoid with subexponential dynamic asymptotic dimension growth is amenable. More generally, every $\sigma$-compact locally compact Hausdorff étale groupoid with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
@@
Jianguo Zhang is partially supported by NSFC 12271165, 12171156, and 12301154
@@ 1. Introduction
The notion of asymptotic dimension was primarily proposed by Gromov in [Gro93] as a large-scale analogue of the Lebesgue covering dimension. The idea gained widespread attention following Yu’s work, demonstrating that the Novikov conjecture holds for groups with finite asymptotic dimension (see [Yu98]). This result motivated numerous mathematicians to explore groups and spaces with finite asymptotic dimension. For instance, Gromov initially established that hyperbolic groups and metric spaces have finite asymptotic dimension (see [Gro93]). Later, Roe provided a more straightforward proof of this finding (see [Roe05]). Following this, Bell and Dranishnikov showed that the free product, the amalgamated free product, and the HNN extension can preserve finite asymptotic dimension (see [BD01]). Four years later, Osin demonstrated that if a finitely generated group $G$ is hyperbolic relative to a collection of subgroups with finite asymptotic dimension, then $G$ itself has finite asymptotic dimension (see [Osi05]). Next, Wright proved that the asymptotic dimension of a finite-dimensional CAT(0) cube complex is bounded above by its dimension (see [Wri12]). Bestvina, Bromberg, and Fujiwara then proved that mapping class groups have finite asymptotic dimension (see [BBF15]). Following that, Behrstock, Hagen, and Sisto proved that all hierarchically hyperbolic groups have finite asymptotic dimension, thereby recovering results related to including mapping class groups, relatively hyperbolic groups, and CAT(0)-cube complexes (see [BHS17]), among many others. For a comprehensive survey of this theory, refer to [BD11]; for recent developments, see [FP21, Tse23, FS24]. Moreover, there are several interactions between this area and $L^{p}$ operator algebras. In the work of [ZZ21], Zhang and Zhou showed that the $L^{p}$ coarse Baum-Connes conjecture holds for proper metric spaces with finite asymptotic dimension when $p\in[1,\infty)$. More recently, Wang, Wang, Zhang, and Zhou demonstrated that any $L^{p}$ Roe algebra for a coarsely uniformly contractible discrete metric space of bounded geometry with finite asymptotic dimension satisfies the persistence approximation property for quantitative $K$-theory of filtered $L^{p}$ operator algebras (see [WWZZ24]).
@@
However, there are still plenty of groups known to have infinite asymptotic dimension, such as the wreath product $\mathbb{Z}\wr\mathbb{Z}$ (see [Dra06]), the first Grigorchuk group (see [Smi07]), Thompson’s group (see [DS11]), Celebrated Gromov’s group in [Gro03] containing an expander. To study the growth of dimension functions of groups with infinite asymptotic dimension, Dranishnikov [Dra06] introduced the notion of asymptotic dimension growth of a metric space. This concept generalizes the notion of finite asymptotic dimension since if the asymptotic dimension function is constant, then the space is of finite asymptotic dimension. Dranishnikov proved that the wreath product of a finitely generated nilpotent group with a finitely generated group of finite asymptotic dimension exhibits polynomial asymptotic dimension growth. Furthermore, he showed that all finitely generated groups have asymptotic dimension growth at most exponential [Dra06]. Additionally, he proved that polynomial asymptotic dimension growth implies Yu’s property A, a weak amenability-type condition introduced by Yu [Yu00]. Combined with Yu’s theorem on the coarse Baum-Connes conjecture [Yu00] implies that spaces with polynomial asymptotic dimension growth satisfy the Baum-Connes conjecture, and hence the Novikov conjecture. Extending Dranishnikov’s result, Ozawa [Oza12] established that a metric space with subexponential asymptotic dimension growth has property A; For another reference, see [Opp14] for the case of discrete metric spaces of bounded geometry.
@@
Many examples of groups and spaces are shown to have subexponential asymptotic dimension growth. An absorbing example is a geodesic coarse median space with finite rank and at most exponential volume growth [ANWZ18], which includes the hyperbolic space and the mapping class group. Another interesting example is the direct product defined as $G=\bigoplus\limits_{n\in\mathbb{N}}f(n)\mathbb{Z}$, where $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ is a strictly increasing function that grows slower than a linear function. When endowed with the direct sum metric, $G$ becomes a discrete metric space of bounded geometry with slow asymptotic dimension growth [CFY09], a specific case of subexponential dimension growth. Besides, Bell examined how various constructions involving groups affect the asymptotic dimension function [Bel05]. To apply this notion to coarse geometry, we will provide a couple of equivalent interlocking definitions of the asymptotic dimension growth in this paper. Moreover, the last equivalent definition lays a good foundation for the theorem in Section 4.
Dynamic asymptotic dimension, developed by Guentner, Willett, and Yu in their paper [GWY17], is a concept of dimension for actions of discrete groups on locally compact spaces, as well as locally compact Hausdorff étale groupoids. This notion is connected to the work of Bartels, Lück, and Reich [BLR08, Bar16] on the Farrel-Jones conjecture in the context of controlled topology and the nuclear dimension in the classification of $C^{*}$-algebras. Subsequently, they extended this idea to dynamical complexity, demonstrating that the Baum-Connes conjecture holds for actions exhibiting finite dynamical complexity [GWY22].
@@
The main aim of this paper is to investigate the growth of dynamic asymptotic dimension function in the context of group actions and groupoids. This study is inspired by Guentner, Willett, and Yu’s work [GWY17], which explores finite dynamic asymptotic dimension in relation to dynamics, topology, coarse geometry, and $C^{*}$-algebras. Their research [GWY22] on dynamical complexity and controlled operator $K$-theory, as well as the work of [ANWZ18] Arzhantseva, Niblo, Wright, and Zhang on characterizing asymptotic dimension growth, forms a solid foundation for our study. These influential works provide valuable methods and results that facilitate our investigation.
@@
In the case of group actions, we will define the dynamic asymptotic dimension growth for actions of discrete groups on compact spaces. This concept can be viewed as a generalization of the finite dynamic asymptotic dimension, and serves as the dynamical version of the asymptotic dimension growth. Analogous to the finite case, we can connect our notion to the work of Bartels et al. [BLR08, Bar16] on the Farrel-Jones conjecture to introduce a new notion called an $h$-BLR action, where $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ is a nondecreasing function. We will show that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. Furthermore, when $h$ grows at a rate slower than $\sqrt{x}$, we derive an equivariant condition related to the above BLR condition, which is formally similar to the definition of asymptotic dimension growth.
@@
In the case of groupoids, we will introduce the dynamic asymptotic dimension growth for locally compact Hausdorff étale groupoids with compact unit space. This notion generalizes the case of group actions. Moreover, we can relate our notion to coarse geometry. More precisely, for a discrete metric space $X$ of bounded geometry, we obtain the following theorem:
@@
This theorem demonstrates that the asymptotic dimension growth for $X$ is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid $G(X)$. The critical point in our proof is to combine the last equivalent definition of asymptotic dimension growth with the proof method that appeared in [[GWY17], Theorem 6.4].
Combining this theorem with Ozawa’s result that the subexponential asymptotic dimension growth implies property A, we derive the following corollary:
@@
This corollary provides an alternative sufficient condition for the amenability of the coarse groupoid. For this result, we can give a concrete example: for a geodesic uniformly locally finite coarse median space with finite rank and at most exponential volume growth $X$, its corresponding coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, and is thus amenable.
With the help of the above BLR condition, we can construct the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions. As a consequence, we have the following theorem:
@@
This theorem is an adaption and a generalization of [[GWY17], Theorem 6.6]. Moreover, it provides a shortcut to construct more examples of group actions with subexponential dynamic asymptotic dimension growth.
The amenable groupoids behave well in several areas, such as the Baum-Connes conjecture [Tu99], the Novikov conjecture [Hig00], and the classification of $C^{*}$-algebras [AR00]. The main implication we explore in this paper is the amenability of groupoids. For this, we develop a key tool that is not only partitions of unity. Using this tool, we can directly prove the following main theorem:
@@
This theorem is a generalization of Corollary 1.2 in the case of slow growth. The idea is also inspired by Guentner, Willett, and Yu’s work [[GWY22], Appendix A ] on finite dynamical complexity for étale groupoids.
The outline of this paper is as follows. In Section 2, we give several equivalent definitions of the asymptotic dimension growth and provide several typical examples. In Section 3, we introduce the dynamic asymptotic dimension growth for group actions and discuss its connection to the work of Bartels, Lück, and Reich. In Section 4, we define the dynamic asymptotic dimension growth for groupoids. Furthermore, we show that the asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid. This allows us to construct the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions. In Section 5, we develop a technical tool more than partitions of unity, which plays a vital role in proving amenability. Finally, in section 6, we find a sufficient condition for amenable groupoids: every $\sigma$-compact, locally compact Hausdorff étale groupoid $\mathcal{G}$ with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
@@ 2. Asymptotic dimension growth for metric spaces
The volume growth of finitely generated groups can be at most exponential. To explore the intermediate dimension growth of such groups, Dranishnikov in [Dra06] proposed the notion of the asymptotic dimension growth, which is a generalization of the finite asymptotic dimension. Additionally, he gave examples of finitely generated groups with polynomial asymptotic dimension growth and showed that all such groups have property A [Dra06]. Later, Ozawa extended this result to the metric spaces with subexponential asymptotic dimension growth [Oza12]. Combined with Yu’s result that the coarse Baum-Connes conjecture holds for spaces with property A, we see that this conjecture holds for spaces with subexponential asymptotic dimension growth. In this sense, subexponential asymptotic dimension growth is weaker than finite asymptotic dimension and stronger than property A. This section will give a couple of equivalent definitions of the asymptotic dimension growth. Moreover, we apply the last equivalent definition to coarse geometry and the proof of Theorem 4.12
@@ 2.1. Equivalent definitions
In this subsection, we will review the notion of the asymptotic dimension growth, and introduce two equivalent definitions. To see this concept more clearly, we will provide several concrete examples.
We review the following notions to define the asymptotic dimension function. Let $(X,d)$ be a metric space and $R>0$. Say that a family $\mathcal{W}$ of subsets of $X$ is $R$-disjoint if $d(W,W^{\prime})>R$ for every $W\neq W^{\prime}$ in $\mathcal{W}$. Say that the family $\mathcal{V}$ is uniformly bounded if $Diam(\mathcal{V})=\sup\{diam(V)\mid V\in\mathcal{V}\}$ is finite. Let $\mathcal{U}$ be a cover of $X$. The Lebesgue number of $\mathcal{U}$, denoted as $L(\mathcal{U})$, is the largest number $R$ such that for every subset $V$ of $X$ with $diam(V)\leq R$ there exists $U\in\mathcal{U}$ such that $V\subset U$. The multiplicity of $\mathcal{U}$, denoted as $m(\mathcal{U})$, is the largest number of elements of $\mathcal{U}$ with nonempty intersection. The $R$-multiplicity of $\mathcal{U}$, denoted as $m_{R}(\mathcal{U})$, is the smallest integer $n$ such that for any $x\in X$, the ball $B(x,R)$ intersects at most $n$ sets of $\mathcal{U}$.
@@
Although the asymptotic dimension growth is not a coarse invariant like the finite asymptotic dimension, it is a quasi-isometric invariant by the following result. Hence, it is a group invariant as any two word metrics on a group are quasi-isometric.
@@
Below, we will introduce three equivalent definitions of asymptotic dimension growth. The first equivalent definition shown in [[ANWZ18], Lemma 2.6] links the previous and the following.
@@
Let $(X,d)$ be a metric space. For any $R>0$, we define $\widetilde{\textnormal{ad}_{X}}(R)$ to be the smallest number over $n\in\mathbb{N}$ for which there exists a uniformly bounded cover $\mathcal{V}=\{V_{j}\}_{j\in J}$ with $m_{R}(\mathcal{V})\leq n+1$. Then $\textnormal{ad}_{X}\approx\widetilde{\textnormal{ad}_{X}}$.
Next, we give the second equivalent definition of the asymptotic dimension growth.
@@
Let $(X,d)$ be a separable metric space. For any $R>0$, we define $\widehat{\textnormal{ad}_{X}}(R)$ to be the smallest number over $n\in\mathbb{N}$ for which there exist uniformly bounded, $R$-disjoint families $\mathcal{W}_{0},\cdots,\mathcal{W}_{n}$ of subsets of $X$ such that $\bigcup\limits_{l=0}^{n}\mathcal{W}_{l}$ is a cover of $X$. Then $\widetilde{\textnormal{ad}_{X}}\approx\widehat{\textnormal{ad}_{X}}$, hence $\textnormal{ad}_{X}\approx\widehat{\textnormal{ad}_{X}}$.
@@
Given $R>0$, $2R>R$. Suppose that there are uniformly bounded, $2R$-disjoint families $\mathcal{W}_{0},\cdots,\mathcal{W}_{\widehat{\textnormal{ad}_{X}}(2R)}$ of subsets of $X$ covering $X$. Put $\mathcal{W}=\bigcup\limits_{l}\mathcal{W}_{l}$, which is obviously a uniformly bounded cover. We claim that $m_{R}(\mathcal{W})\leq\widehat{\textnormal{ad}_{X}}(2R)$. By definition of $R$-multiplicity, for any $x\in X$, there are at most $\widehat{\textnormal{ad}_{X}}(2R)+1$ of sets of $\mathcal{W}$ having non-empty intersection with the closed ball $B(x,R)$. Otherwise, there must be two sets $W$ and $W^{\prime}$ from the same family $\mathcal{W}_{i}$ for some $i$ such that $W\cap B(x,R)\neq\varnothing$ and $W^{\prime}\cap B(x,R)\neq\varnothing$, since there are only $\widehat{\textnormal{ad}_{X}}(2R)+1$ different families. Then $d(W,W^{\prime})\leq 2R$, which contradicts that the family $\mathcal{W}_{i}$ is $2R$-disjoint. Hence, $m_{R}(\mathcal{W})\leq\widehat{\textnormal{ad}_{X}}(2R)$ as claimed, and thus yields $\widetilde{\textnormal{ad}_{X}}\preceq\widehat{\textnormal{ad}_{X}}$.
Conversely, given $R>0$, suppose $\mathcal{V}=\{V_{j}\}_{j\in J}$ is a unifromly bounded cover with $m_{R}(\mathcal{V})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$. Since $X$ is separable, $X$ is second countable, so a cover of $X$ must have a countable subcover. Thus we may assume that the cover $\mathcal{V}$ is countable. We first claim that there exist uniformly bounded, $R$-disjoint families $\mathcal{V}_{0},\cdots,\mathcal{V}_{\widetilde{\textnormal{ad}_{X}}(R)}$ of subsets of $X$ covering $X$. We prove the first claim in three steps.

### v2

@@ Abstract.
We introduce the notion of dynamic asymptotic dimension growth for actions of discrete groups on compact spaces, and more generally for locally compact étale groupoids. Using the work of Bartels, Lück, and Reich, we bridge asymptotic dimension growth for countable discrete groups with our notion for their group actions, thereby providing numerous concrete examples. Moreover, we demonstrate that the asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid. Consequently, we deduce that the coarse groupoid with subexponential dynamic asymptotic dimension growth is amenable. More generally, we show that every $\sigma$-compact locally compact Hausdorff étale groupoid with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
@@
Jianguo Zhang is supported by NSFC 12271165, 12171156, and 12301154
Dapeng Zhou is supported by NSFC 12271165
@@ 1. Introduction
The notion of asymptotic dimension was primarily proposed by Gromov in [Gro93] as a large-scale analogue of the Lebesgue covering dimension. This idea gained widespread attention following Yu’s work, demonstrating that the Novikov conjecture holds for groups with finite asymptotic dimension (see [Yu98]). This pivotal result motivated numerous mathematicians to investigate groups and spaces with finite asymptotic dimension. For instance, Gromov initially established that hyperbolic groups and metric spaces have finite asymptotic dimension (see [Gro93]). Subsequently, Roe provided a more straightforward proof of this finding (see [Roe05]). Building on this, Bell and Dranishnikov showed that the free product, the amalgamated free product, and the HNN extension can preserve finite asymptotic dimension (see [BD01]). Four years later, Osin demonstrated that if a finitely generated group $G$ is hyperbolic relative to a collection of subgroups with finite asymptotic dimension, then $G$ itself has finite asymptotic dimension (see [Osi05]). Further advancements were made by Wright, who proved that the asymptotic dimension of a finite-dimensional CAT(0) cube complex is bounded above by its dimension (see [Wri12]). Additionally, Bestvina, Bromberg, and Fujiwara established that mapping class groups have finite asymptotic dimension (see [BBF15]). Following this, Behrstock, Hagen, and Sisto proved that all hierarchically hyperbolic groups have finite asymptotic dimension, thereby recovering results including mapping class groups, relatively hyperbolic groups, and CAT(0)-cube complexes (see [BHS17]), among many others. For a comprehensive survey of this theory, refer to [BD11]; for recent developments, see [FP21, Tse23, FS24].
Moreover, finite asymptotic dimension has found applications in the domain of $L^{p}$ operator algebras as well. Analogous to the coarse Baum-Connes conjecture, the third and fourth authors established the $L^{p}$ coarse Baum-Connes conjecture for $p\in[1,\infty)$ and proved that this conjecture holds for proper metric spaces with finite asymptotic dimension (see [ZZ21]). Subsequently, the current authors provided an alternative proof of this result through the framework of $C_{0}$ coarse geometry (see [WWZZ23]). Recently, the authors investigated the persistence approximation property regarding quantitative $K$-theory of filtered $L^{p}$ operator algebras and demonstrated that any $L^{p}$ Roe algebra for a coarsely uniformly contractible discrete metric space with bounded geometry and finite asymptotic dimension satisfies the persistence approximation property (see [WWZZ24]).
@@
However, plenty of groups are known to have infinite asymptotic dimension, including the wreath product $\mathbb{Z}\wr\mathbb{Z}$ (see [Dra06]), the first Grigorchuk group (see [Smi07]), Thompson’s group (see [DS11]),and the Celebrated Gromov’s group in [Gro03] containing an expander. To study the growth of dimension functions of groups with infinite asymptotic dimension, Dranishnikov [Dra06] introduced the notion of asymptotic dimension growth for a metric space. This concept extends the notion of finite asymptotic dimension, since a constant asymptotic dimension function indicates finite asymptotic dimension. Dranishnikov also established that the wreath product of a finitely generated nilpotent group with a finitely generated group of finite asymptotic dimension exhibits polynomial asymptotic dimension growth. Furthermore, he proved that polynomial asymptotic dimension growth implies Yu’s property A [Dra06], a weak amenability-type condition introduced by Yu [Yu00]. In conjunction with Yu’s theorem on the coarse Baum-Connes conjecture [Yu00], it follows that spaces with polynomial asymptotic dimension growth satisfy the coarse Baum-Connes conjecture, and consequently, the Novikov conjecture. Building on Dranishnikov’s insights, Ozawa showed that a metric space with subexponential asymptotic dimension growth has property A [Oza12]; For additional reference, see [Opp14] regarding discrete metric spaces of bounded geometry. In this context, subexponential asymptotic dimension growth is considered weaker than finite asymptotic dimension but stronger than property A.
@@
Many examples of groups and spaces exhibit subexponential asymptotic dimension growth. A fascinating example is a geodesic coarse median space with finite rank and at most exponential volume growth that has subexponential asymptotic dimension growth [ANWZ18]. This space includes the hyperbolic space and the mapping class group. Another intriguing example is the direct product $G=\bigoplus\limits_{n\in\mathbb{N}}f(n)\mathbb{Z}$, where $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ is a strictly increasing function that grows slower than a linear function. When endowed with the direct sum metric, $G$ becomes a discrete metric space of bounded geometry with slow asymptotic dimension growth [CFY09], a special case of subexponential asymptotic dimension growth. Additionally, Bell examined how various constructions involving groups affect the asymptotic dimension function [Bel05]. In this paper, we will introduce three equivalent interlocking definitions of asymptotic dimension growth. The first equivalent definition is derived from [[ANWZ18], Lemma 2.6]; the third equivalent definition uses the language of coarse geometry and will be applied to the proof of Theorem 1.1.
Inspired by Gromov’s theory of asymptotic dimension, Guentner, Willett, and Yu introduced the dynamic asymptotic dimension in their seminal paper [GWY17], which is a concept of dimension for actions of discrete groups on locally compact spaces, as well as locally compact Hausdorff étale groupoids. Since its inception, this notion has various applications. It provides upper bounds on the nuclear dimension of the resulting groupoid $C^{*}$-algebras, as demonstrated in [GWY17, DS18, ALSS21, HW24], and even extends to twisted groupoid $C^{*}$-algebras [CDGHV24, BL24]. Additionally, it establishes a sufficient condition for Matui’s HK conjecture in low dimensions [BDGW23]. Furthermore, the dynamic asymptotic dimension is intimately connected to other key concepts in the field. It correlates closely with the diagonal dimension of sub-$C^{*}$-algebras [LLW23], Kerr’s tower dimension [Ker20], and the conditions formulated by Bartels, Lück, and Reich [BLR08, Bar16]. In a recent study [Bon24], Bönicke enriched the understanding of this dimension by establishing several permanence properties including estimates for products and unions of groupoids, as well as comparing the asymptotic dimension of the resulting coarse space with the dynamic asymptotic dimension of the underlying groupoid.
@@
However, group actions and groupoids with infinite dynamic asymptotic dimension have not been extensively explored. The aim of this paper is to contribute to their study by examining the growth of the dynamic asymptotic dimension function for these group actions and groupoids. This study is inspired by the work of Guentner, Willett, and Yu in [GWY17] regarding finite dynamic asymptotic dimension for group actions and étale groupoids, along with their subsequent research in [GWY24] verifying the Baum-Connes conjecture for actions with finite dynamical complexity. It should be emphasized that when the dynamic asymptotic dimension is either zero or one, the finite dynamic asymptotic dimension implies finite dynamical complexity. The work of Arzhantseva, Niblo, Wright, and Zhang in [ANWZ18] on characterizing asymptotic dimension growth forms a solid foundation for our study. These influential works provide valuable methods and results that facilitate our investigation.
@@
In the case of group actions, we will introduce the concept of dynamic asymptotic dimension growth for actions by discrete groups on compact Hausdorff spaces in Definition 3.2. This notion can be viewed as the dynamical counterpart of asymptotic dimension growth in an appropriate sense, and it serves as a generalization of finite dynamic asymptotic dimension. The dynamic asymptotic dimension has been estimated for various group actions [ALSS21, CJMST23, GWY17]. Building upon these techniques and findings, we will provide examples of group actions exhibiting dynamic asymptotic dimension growth. For this, we link our notion to the works of Bartels, Lück, and Reich on the Farrel-Jones conjecture [BLR08, Bar16] by introducing the $h$-BLR action, where $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ is a nondecreasing function. Furthermore, when $h$ grows at a rate slower than $\sqrt{x}$, we derive an alternative equivalent definition of the $h$-BLR action, which is formally similar to the definition of asymptotic dimension growth. Additionally, we will demonstrate that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. By applying this result within the domain of coarse geometry, we will bridge the asymptotic dimension growth for countable discrete groups with the dynamic asymptotic dimension growth for their actions, as elaborated in Section 4. This allows us to present several representative examples of group actions with subexponential dynamic asymptotic dimension growth.
@@
In the case of étale groupoids, we will define the concept of dynamic asymptotic dimension growth for locally compact Hausdorff étale groupoids, thereby extending our earlier notion for actions of discrete groups on compact spaces. Initially, we define the dynamic asymptotic dimension growth of an étale groupoid with compact unit space in Definition 4.6. For a groupoid with non-compact unit space, our definition reduces to the compact case via the Alexandrov groupoid, as outlined in[[KKLRU21], Definition 7.7]. Within the framework of coarse geometry, we establish a connection between our concept and the asymptotic dimension growth, leading to the following theorem:
@@
This theorem is our principal motivation, as it is related to the coarse Baum-Connes conjecture. It establishes that the asymptotic dimension growth of $X$ is equivalent to the dynamic asymptotic dimension growth of its associated coarse groupoid $G(X)$, which generalizes [[GWY17], Theorem 6.4]. The pivotal step in our proof involves integrating the third equivalent definition of asymptotic dimension growth with the proof technique presented in [[GWY17], Theorem 6.4].
Combining this theorem with Ozawa’s result that subexponential asymptotic dimension growth implies property A [Oza12], we derive the following corollary:
@@
This corollary offers an alternative sufficient condition for the amenability of the coarse groupoid. To illustrate this result, we can present a concrete example: for a geodesic uniformly locally finite coarse median space $X$ with finite rank and at most exponential volume growth, its corresponding coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, thereby confirming its amenability.
Combining Theorem 1.1, Ozawa’s result [[Oza12], Theorem 1] and the fact that an $h$-BLR action has dynamic asymptotic dimension function at most $h$, we can establish the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions, as elaborated in Theorem 4.21. Consequently, we arrive at the following theorem:
@@
This theorem is an adaption and a generalization of [[GWY17], Theorem 6.6]. Moreover, it provides a shortcut to construct more concrete examples of group actions with subexponential dynamic asymptotic dimension growth. For example, the wreath product $\mathbb{Z}\wr\mathbb{Z}$ has infinite asymptotic dimension and polynomial asymptotic dimension growth. According to this theorem, $\mathbb{Z}\wr\mathbb{Z}$ admits a free, minimal action on the Cantor set, and the dynamic asymptotic dimension of this action is infinite and grows polynomially.
The amenable groupoids has several major results in various aspects, such as the Baum-Connes conjecture [Tu99], the Novikov conjecture [Hig00], and the classification of $C^{*}$-algebras [AR00]. The main implication we explore in this paper is the amenability of groupoids. To this end, we develop a key tool known as generalized partitions of unity. Using this tool, we can directly establish the following main theorem:
@@
This theorem is a generalization of Corollary 1.2 in the case of slow growth. Our proof is more closely modeled on the beautiful approach by Guentner, Willett, and Yu’s work [[GWY24], Appendix A ] on finite dynamical complexity for étale groupoids. Using a similar approach, we can prove that every locally compact Hausdorff étale groupoid with finite dynamic asymptotic dimension is amenable, which strengthens the result of [GWY17] by removing the freeness assumption for groupoids.
The outline of this paper is as follows. In Section 2, we present three equivalent definitions of the asymptotic dimension growth and illustrate them with four typical examples. In Section 3, we introduce the concept of dynamic asymptotic dimension growth for group actions, discussing its connection to the work of Bartels, Lück, and Reich. In Section 4, we define the dynamic asymptotic dimension growth for étale groupoids. Furthermore, we demonstrate that the asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid. This equivalence allows us to establish a connection between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for their group actions. In Section 5, we develop a technical tool known as a generalized partition of unity, which is crucial for proving the amenability of groupoids. Finally, in section 6, we identify a sufficient condition for amenable groupoids: every $\sigma$-compact, locally compact Hausdorff étale groupoid $\mathcal{G}$ with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
@@ 2. Asymptotic dimension growth for metric spaces
The volume growth of finitely generated groups can be at most exponential. To explore the intermediate dimension growth of these groups, Dranishnikov proposed the concept of asymptotic dimension growth in [Dra06], which generalizes the notion of finite asymptotic dimension. In this section, we will introduce three equivalent definitions of asymptotic dimension growth. The third equivalent definition employs the terminology of coarse geometry, which can be used in the proof of Theorem 4.16 in Section 4. To facilitate a more intuitive understanding of this concept, we provide examples of spaces exhibiting exponential growth, subexponential growth, polynomial growth, and slow growth.
@@ 2.1. Equivalent definitions
In this subsection, we will review the concept of asymptotic dimension growth and present two equivalent definitions. To enhance clarity, we will illustrate this concept with several concrete examples.
To clarify the definition of asymptotic dimension function, we will outline some elementary concepts. Let $(X,d)$ be a metric space and $R>0$. A family $\mathcal{W}$ of subsets of $X$ is termed $R$-disjoint if the distance $d(W,W^{\prime})>R$ for every pair of distinct subsets $W\neq W^{\prime}$ within $\mathcal{W}$. A family $\mathcal{V}$ is termed uniformly bounded if its diameter, defined as $Diam(\mathcal{V})=\sup\{diam(V)\mid V\in\mathcal{V}\}$, is finite.
Let $\mathcal{U}$ be a cover of $X$. The Lebesgue number of $\mathcal{U}$, denoted by $L(\mathcal{U})$, is the maximum number $R$ such that for every subset $V$ of $X$ with diameter $diam(V)\leq R$, there exists a set $U\in\mathcal{U}$ such that $V\subseteq U$. The multiplicity of $\mathcal{U}$, denoted as $m(\mathcal{U})$, is the largest number of elements of $\mathcal{U}$ with nonempty intersection. The $R$-multiplicity of $\mathcal{U}$, denoted by $m_{R}(\mathcal{U})$, is the smallest integer $n$ such that for any $x\in X$, the ball $B(x,R)$ intersects at most $n$ sets of $\mathcal{U}$.
With this in mind we can recall the definition of asymptotic dimension function.
@@
Although the asymptotic dimension growth is not a coarse invariant like the finite asymptotic dimension, it turns out to be a quasi-isometric invariant by the following result. Thus, it is a group invariant because any two word metrics on a group are quasi-isometric.
@@
Below, we will present three equivalent definitions of asymptotic dimension growth. The first of these, as demonstrated in [[ANWZ18], Lemma 2.6], serves as a bridge between the preceding and subsequent definitions.
@@
Let $(X,d)$ be a metric space. For any $R>0$, we define $\widetilde{\textnormal{ad}_{X}}(R)$ to be the smallest number over $n\in\mathbb{N}$ for which there exists a uniformly bounded cover $\mathcal{V}=\{V_{j}\}_{j\in J}$ with $m_{R}(\mathcal{V})\leq n+1$. Then $\widetilde{\textnormal{ad}_{X}}\approx\textnormal{ad}_{X}$.
The succeeding proposition offers the second equivalent definition of asymptotic dimension growth. This pivotal step is essential in establishing the third equivalent definition.
@@
Let $(X,d)$ be a separable metric space. For any $R>0$, we define $\widehat{\textnormal{ad}_{X}}(R)$ to be the smallest number over $n\in\mathbb{N}$ for which there exist uniformly bounded, $R$-disjoint families $\mathcal{W}_{0},\cdots,\mathcal{W}_{n}$ of subsets of $X$ such that $\bigcup\limits_{l=0}^{n}\mathcal{W}_{l}$ is a cover of $X$. Then $\widehat{\textnormal{ad}_{X}}\approx\widetilde{\textnormal{ad}_{X}}$, hence $\widehat{\textnormal{ad}_{X}}\approx\textnormal{ad}_{X}$.
@@
Given $R>0$, it follows that $2R>R$. Suppose there exist uniformly bounded, $2R$-disjoint families $\mathcal{W}_{0},\cdots,\mathcal{W}_{\widehat{\textnormal{ad}_{X}}(2R)}$ of subsets of $X$ that cover $X$. Put $\mathcal{W}=\bigcup\limits_{l}\mathcal{W}_{l}$, which is obviously a uniformly bounded cover. We claim that $m_{R}(\mathcal{W})\leq\widehat{\textnormal{ad}_{X}}(2R)$. By the definition of $R$-multiplicity, for any $x\in X$, there are at most $\widehat{\textnormal{ad}_{X}}(2R)+1$ sets in $\mathcal{W}$ having non-empty intersection with the closed ball $B(x,R)$. Otherwise, there must be two sets $W$ and $W^{\prime}$ from the same family $\mathcal{W}_{i}$ for some $i$ such that both $W\cap B(x,R)\neq\varnothing$ and $W^{\prime}\cap B(x,R)\neq\varnothing$, since there are only $\widehat{\textnormal{ad}_{X}}(2R)+1$ different families. This implies that $d(W,W^{\prime})\leq 2R$, contradicting the $2R$-disjoint property of the family $\mathcal{W}_{i}$. Therefore, $m_{R}(\mathcal{W})\leq\widehat{\textnormal{ad}_{X}}(2R)$ as claimed, and thus yields $\widetilde{\textnormal{ad}_{X}}\preceq\widehat{\textnormal{ad}_{X}}$.
Conversely, given $R>0$, suppose $\mathcal{V}=\{V_{j}\}_{j\in J}$ is a unifromly bounded cover with $m_{R}(\mathcal{V})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$. Since $X$ is separable, $X$ is second countable, implying that any cover of $X$ must have a countable subcover. Therefore, we can assume that the cover $\mathcal{V}$ is countable. We initially claim that there exist uniformly bounded, $R$-disjoint families $\mathcal{V}_{0},\cdots,\mathcal{V}_{\widetilde{\textnormal{ad}_{X}}(R)}$ of subsets of $X$ covering $X$. This claim will be demonstrated through a three-step process.